OpenStudy (anonymous):

3 consecutive odd ontegers are such that the square of the third integer is 65 less than the sum of the squares of the first two. one solution is -7, -5, -3. find 3 other consecutive odd integers that also satisfy the given conditions

OpenStudy (anonymous):

11, 13, 15 :))

OpenStudy (anonymous):

OMG your like super smart cud u explain y

OpenStudy (anonymous):

(x+4)^2=x^2+(x+2)^2-65

OpenStudy (anonymous):

x^2-4x-77=0 (x-11)(x+7)=0 x=11 or x=-7

OpenStudy (anonymous):

-7, -5, -3 11,13,15

OpenStudy (anonymous):

our numbers are: x -------->1st number x+2 ----->2nd number x+4-------> 3rd number and our equation is: x^2 + (x+2)^2 - 65 = (x+4)^2 then just simplify: it will be: x^2 - 4x -77 = 0 then get the value of x wich is 11, so: x = 11 x+2 = 13 X+4 = 15 :))

OpenStudy (anonymous):

omg thank you